# RealEngine: Simulating Autonomous Driving in Realistic Context

### v1

@@ 1 Introduction
Autonomous driving (AD) methods yin2021center ; li2022homogeneous ; ye2023fusionad ; hu2023planning ; wen2023dilu ; fu2024drive have advanced rapidly, largely due to the introduction of diverse driving datasets nuplan ; caesar2020nuscenes ; montali2024waymo and simulators dosovitskiy2017carla . These resources facilitate research by supporting model training, testing, and evaluation across a variety of virtual and real driving environments, utilizing a wealth of multimodal sensor, including cameras and LiDAR.
@@
However, existing AD datasets and simulators have fundamental limitations that prevent them from fully capturing real-world driving scenarios and challenges, diminishing their credibility and usefulness in practical applications. For instance, real driving data typically provides only pre-existing driving trajectories nuplan ; caesar2020nuscenes , allowing for open-loop evaluation without immediate feedback or interaction with the trajectories planned by driving agents. This results in a significant discrepancy between simulated and actual driving behavior. The static nature of recorded datasets, where other vehicles do not react to the actions of the ego vehicle, creates a scenario fundamentally different from real-world situations. Although the closed-loop simulator CARLA dosovitskiy2017carla addresses this issue by providing real-time feedback to driving agents, it relies on manual 3D modeling and graphics engines, which lacks realism and exhibits a substantial appearance gap from actual driving scenarios. Consequently, models trained in this environment may struggle to handle real-world driving conditions effectively. Further, previous benchmarks primarily focus on non-collision and normal driving scenarios, which limits the models’ ability to address unseen risks encountered in the real world.
@@
To advance the field, we introduce RealEngine, a pioneering autonomous driving simulation platform capable of rendering realistic multimodal sensor data efficiently and supporting closed-loop simulation. It is distinguished by the following features: (i) Realistic scene rendering, closely resembling the real world to minimize domain discrepancies between simulated and actual driving environments, allowing both camera images and LiDAR point clouds; (ii) Closed-loop simulation, enabling driving along free-form trajectories planned by agents while providing corresponding feedback; (iii) Supporting diverse scenarios, including a wide range of hazardous driving situations to facilitate comprehensive safety-critical evaluations of driving agents; (iv) Multi-agent co-operation and interaction, closely approximating various real-world conditions in terms of driving dynamics and scene complexity. As summarized in Table 1, no existing benchmarks or simulators meet these fundamental requirements simultaneously, as well as a spectrum of driving simulation focused features.
@@
Our RealEngine is founded on the innovative concept of seamlessly integrating scene reconstruction with the composition of traffic participants. We reconstruct the background scene and foreground traffic participants separately with the corresponding real-world data. This approach simultaneously addresses key limitations of existing scene reconstruction methods chen2023periodic ; wu2023mars ; yang2023emernerf ; yan2024street ; chen2024omnire ; turki2023suds : (i) Unable to acquire occluded regions within a scene; (ii) Inferior performance in synthesizing novel viewpoints of traffic participants; (iii) Limited choice of possible traffic participants. Our decomposed design naturally supports flexible scene editing while enabling the concurrent operation of multiple driving agents, as required for realistic driving simulations. By merging the background scene with diverse traffic participants, we can efficiently simulate a wide range of high-quality, unique driving scenarios tailored to any specific requirements. We generate a diverse range of driving scenarios and simulate three driving categories: non-reactive, safety test, and multi-agent interaction. This offers a high quality, reliable, flexible, and comprehensive benchmark for assessing the real-world performance of driving agents.
@@
DATA. CitySim robinson2009citysim / NGSIM usdot2016ngsim
Bench2Drive jia2024bench2drive
nuPlan nuplan / Navsim Dauner2024NEURIPS
nuScenes caesar2020nuscenes / Waymo dataset sun2020scalability
GEN. MagicDrive gao2023magicdrive / DriveDreamer wang2023drivedreamer
SimGen zhou2024simgen
W.M. KiGRAS zhao2024kigras / SMART wu2024smart
MUVO bogdoll2023muvo
Vista gao2024vista / GAIA-1 hu2023gaia
SIM. Waymax gulino2024waymax
SUMO krajzewicz2012recent / LimSim wenl2023limsim
CARLA dosovitskiy2017carla
STRIVE rempe2022generating
MetaDrive li2022metadrive
Unisim yang2023unisim / OAsim yan2024oasim
NeuroNCAP ljungbergh2024neuroncap
DriveArena yang2024drivearena
HUGSIM zhou2024hugsim
@@ 2 Related work
Autonomous driving. Recent research in autonomous driving have shifted from addressing individual tasks to exploring end-to-end planning, enabling the progressive execution from perception to ego-planning within a unified framework. Early works casas2021mp3 ; hu2022stp3 ; Chitta2023PAMI implement this by leveraging simplified auxiliary tasks, which limits the final performance. In contrast, UniAD hu2023planning and VAD jiang2023vad have advanced this paradigm by integrating a broader spectrum of driving tasks, achieving notable progress in planning tasks by producing explicit intermediate results. Additionally, recent approaches present sparse representations sparsead ; sparsedrive , employ lightweight diffusion-based generation liao2024diffusiondrive to efficiently facilitate end-to-end planning. For enabling these driving methods to be evaluated extensively and authentically, we provide a reliable testing platform here.
@@
Sensor simulation. The geometric reconstruction of extensive urban spaces, such as streets and cities, has played a pivotal role in autonomous driving sun2020scalability ; geiger2012we ; caesar2020nuscenes . Recently, Gaussian splatting kerbl3Dgaussians based methods have been introduced to model dynamic urban scenes. PVG chen2023periodic incorporates periodic vibration at each Gaussian primitive to represent static and dynamic objects uniformly. At the same time, explicitly decomposing scenes into independent entities has become common practice, as seen in works such as StreetGaussians yan2024street , DrivingGaussians zhou2024drivinggaussian , HUGS hugs and OmniRe chen2024omnire . Recently, LiDAR simulation studies have focused on using real data for improved realism. For instance, LiDARsim manivasagam2020lidarsim and PCGen li2023pcgen use multi-step, data-driven methods to simulate point clouds from real data. Additionally, works such as tao2023lidarnerf ; zhang2024nerf ; zheng2024lidar4d ; xue2024geonlf ; tao2024alignmif ; wu2024dynamic leverage NeRF mildenhall2020nerf for scene reconstruction and LiDAR simulation. Recent works jiang2025gslidar ; zhou2024lidarrt introduce Gaussian splatting kerbl3Dgaussians into the LiDAR reconstruction task, achieving improved reconstruction quality and rendering speed. However, sensor-acquired data is prone to occlusions, leading to information loss within scenes. Additionally, reconstructed dynamic objects can distort when viewed from alternative perspectives, limiting scene editability and novel view synthesis. To address these challenges, we propose modeling the foreground and background separately allowing them to be composed in a flexible manner, resulting in a comprehensive driving simulation platform.
@@
Closed-loop simulation in realistic settings. Closed-loop simulation gulino2024waymax ; nuplan ; dosovitskiy2017carla ; li2022metadrive is crucial for the evaluation and deployment of AD planning systems, collecting sufficient driving statistics and improving the emergency response capability of driving agents. To make simulators more realistic, recent research has explored using existing real-world driving datasets. Bench2Drive jia2024bench2drive improves the CARLA benchmark by reconstructing the data format to be more aligned with the nuScenes dataset, bridging the gap between reinforcement learning planners and end-to-end planners. Navsim Dauner2024NEURIPS imitates closed-loop evaluation to adjust the nuPlan benchmark, introducing more comprehensive and practical metrics. With the emergence of novel 3D rendering and generation research, some works integrate these techniques into existing datasets, better simulating and constructing richer and more diverse scenarios. NeuroNCAP ljungbergh2024neuroncap utilizes NeRF mildenhall2020nerf to render novel surrounding views and creates collision scenarios to enhance measurement. Relying on powerful generative models and extensive driving data, DriveArena yang2024drivearena is capable to yield controllable and abundant real-world scenarios and realize closed-loop simulation. However, these simulators all fail to meet the fundamental criteria as mentioned earlier. This motivates the introduction of our RealEngine, a more reliable and powerful benchmark for assessing the real-world performance of driving agents.
@@ 3 Method
We present RealEngine, a reliable and comprehensive autonomous driving simulation platform capable of rendering realistic multimodal sensor data efficiently and supporting closed-loop simulation. It is composed of: (i) Simulator infrastructure including background scene and traffic participants as well as their compositions; (ii) Closed-loop driving simulation; (iii) Assessment of driving agents. The working mechanism of RealEngine is depicted in Figure 1.
@@ 3.1 Simulator infrastructure
The advancement of recent generative models has been recently exploited for closed-loop driving simulation. For example, DriveArena yang2024drivearena streamlines a couple of generative functions, such as interactive traffic flows and novel scene synthesis, allowing to simulate virtual driving scenes. However, this approach suffers from several key drawbacks: (1) Need for large manually labelled training data for video generative model optimization; (2) Substantial running overhead, even without support to multimodal sensor data such as LiDAR point clouds (3) Inability to support multi-agent co-operation and interaction. (4) Dependence on a pre-defined high-definite map; (5) Low controllability on the traffic scenario including background scene and traffic participants, consequently causing frequent situational inconsistency over space and time; (6) Low spatiotemporal realism. While some of the above issues, such as realistic scene rendering, more flexible traffic scenario editing, and no need for training data collection, can be addressed by NeuroNCAP ljungbergh2024neuroncap , it narrowly focuses on safety test of driving agents. Further, it cannot well generalize to free form trajectories otherwise the novel view synthesis will degraded substantially.
@@ 3.1.1 Background scene models
To reconstruct realistic background environments, we adopt StreetGaussians yan2024street for camera images and GS-LiDAR jiang2025gslidar for LiDAR point clouds, chosen for their high rendering efficiency and cross-modal fidelity. Importantly, RealEngine maintains flexibility—alternative methods such as PVG chen2023periodic , OmniRe chen2024omnire , and LiDAR-RT zhou2024lidarrt can be seamlessly incorporated into our system as modular components.
Pose calibrations in nuPlan nuplan is relatively coarse, posing challenges for accurate scene reconstruction. A widely used conventional method for camera pose correction yan2024street ; chen2024omnire ; hess2024splatad involves learning a trainable correction matrix that adjusts the camera pose automatically during training. However, this method is only effective for small pose deviations and fails to converge on nuPlan, leading to suboptimal results. To address this issue, we propose a pose correction method based on LiDAR point cloud registration.
As the geometric information of the background scene remains fixed in the world coordinate system, we can align the ego vehicle’s poses across different frames by registering LiDAR points transformed to the world. We remove dynamic vehicles from the LiDAR data based on annotation information and filter out ground points which vary significantly across frames. All remaining point clouds are transformed into the world coordinate system and truncated within a predefined region to ensure consistency across frames. We select the central LiDAR frame ${\bm{P}}_{\mathrm{refer}}$ as the reference frame and apply a learnable correction matrix ${\bm{M}}$ to transform the LiDAR frames ${\bm{P}}$. The transformed LiDAR frames are then compared with the reference frame using the Chamfer Distance (C-D) fan2017chamfer as the loss function:
@@
Additionally, we leverage learnable image exposure transformations to handle cross-camera appearance variations, and utilizes video generative prior yang2024drivex ; yu2024viewcrafter to optimize the scene model across multiple trajectories in Section A.2. For further details on camera image and LiDAR point cloud reconstruction, please refer to Appendix A. By leveraging these advanced reconstruction techniques, we achieve precise multi-modal background reconstruction with high rendering efficiency, enabling flexible scene editing and realistic simulation of autonomous driving agents.
@@ 3.1.2 Traffic participants models
Traffic participants (e.g., vehicles, bicycles) are essential for realistic driving scenarios. To enable high-fidelity and diverse simulations, we curate a comprehensive set of 3D participant models. Conventional approaches yan2024street ; chen2024omnire ; zhou2024hugsim insert reconstructed Gaussian splatting primitives into the background scene according to predefined trajectories. While providing high-quality rendering from training viewpoints, they degrade significantly when rendering resolution, viewpoint, or object distance changes yan2024multi ; Yu2024MipSplatting ; sags . This degradation causes visual artifacts, especially in close-range interactions (e.g., collisions), and also introduces inconsistencies in lighting and shadows between foreground objects and the background scene.
To address these issues, RealEngine represents traffic participants using 3D meshes, ensuring consistent geometry across all viewpoints and distances. To achieve seamless integration, we employ a diffusion model to guide learning of scene-consistent lighting and shading, as further detailed in Section 3.1.3. Meshes are rendered using ray tracing to produce both RGB images and depth maps, enabling accurate occlusion reasoning between participants and scenes.
The 3D meshes are sourced from two complementary pipelines: (i) High-quality meshes manually created for key traffic objects. (ii) Meshes reconstructed from real-world datasets, including 360-degree images from CO3D reizenstein21co3d and 3DRealCar du20243drealcar . These are processed via 3D Gaussian Splatting (3DGS) for reconstruction, followed by mesh extraction for flexible use in scene composition.

### v2

@@ 1 Introduction
Autonomous driving (AD) methods (Yin et al., 2021; Li et al., 2022b; Ye et al., 2023; Hu et al., 2023b; Wen et al., 2023; Fu et al., 2024) have advanced rapidly, largely due to the introduction of diverse driving datasets (H. Caesar, 2021; Caesar et al., 2020; Montali et al., 2024) and simulators (Dosovitskiy et al., 2017). These resources facilitate research by supporting model training, testing, and evaluation across a variety of virtual and real driving environments, utilizing a wealth of multimodal sensor, including cameras and LiDAR.
@@
However, existing AD datasets and simulators have fundamental limitations that prevent them from fully capturing real-world driving scenarios and challenges, diminishing their credibility and usefulness in practical applications. For instance, real driving data typically provides only pre-existing driving trajectories (H. Caesar, 2021; Caesar et al., 2020), allowing for open-loop evaluation without immediate feedback or interaction with the trajectories planned by driving agents. This results in a significant discrepancy between simulated and actual driving behavior. The static nature of recorded datasets, where other vehicles do not react to the actions of the ego vehicle, creates a scenario fundamentally different from real-world situations. Although the closed-loop simulator CARLA (Dosovitskiy et al., 2017) addresses this issue by providing real-time feedback to driving agents, it relies on manual 3D modeling and graphics engines, which lacks realism and exhibits a substantial appearance gap from actual driving scenarios. Consequently, models trained in this environment may struggle to handle real-world driving conditions effectively. Further, previous benchmarks primarily focus on non-collision and normal driving scenarios, which limits the models’ ability to address unseen risks encountered in the real world.
@@
To advance the field, we introduce RealEngine, a pioneering autonomous driving simulation platform capable of rendering realistic multimodal sensor data efficiently and supporting closed-loop simulation. It is distinguished by the following features: (i) Realistic scene rendering, closely resembling the real world to minimize domain discrepancies between simulated and actual driving environments, allowing both camera images and LiDAR point clouds; (ii) Closed-loop simulation, enabling driving along free-form trajectories planned by agents while providing corresponding feedback; (iii) Supporting diverse scenarios, including a wide range of hazardous driving situations to facilitate comprehensive safety-critical evaluations of driving agents; (iv) Multi-agent co-operation and interaction, closely approximating various real-world conditions in terms of driving dynamics and scene complexity. As summarized in Table˜1, no existing benchmarks or simulators meet these fundamental requirements simultaneously, as well as a spectrum of driving simulation focused features.
@@
Our RealEngine is founded on the innovative concept of seamlessly integrating scene reconstruction with the composition of traffic participants. We reconstruct the background scene and foreground traffic participants separately with the corresponding real-world data. This approach simultaneously addresses key limitations of existing scene reconstruction methods (Chen et al., 2023; Wu et al., 2023; Yang et al., 2023a; Yan et al., 2024b; Chen et al., 2024; Turki et al., 2023): (i) Unable to acquire occluded regions within a scene; (ii) Inferior performance in synthesizing novel viewpoints of traffic participants; (iii) Limited choice of possible traffic participants. Our decomposed design naturally supports flexible scene editing while enabling the concurrent operation of multiple driving agents, as required for realistic driving simulations. By merging the background scene with diverse traffic participants, we can efficiently simulate a wide range of high-quality, unique driving scenarios tailored to any specific requirements. We generate a diverse range of driving scenarios and simulate three driving categories: non-reactive, safety test, and multi-agent interaction. This offers a high quality, reliable, flexible, and comprehensive benchmark for assessing the real-world performance of driving agents.
@@
DATA. CitySim (Robinson et al., 2009)
Bench2Drive (Jia et al., 2024)
nuPlan (H. Caesar, 2021) / Navsim (Dauner et al., 2024)
nuScenes (Caesar et al., 2020) / Waymo dataset (Sun et al., 2020)
GEN. MagicDrive (Gao et al., 2023) / DriveDreamer (Wang et al., 2023a)
SimGen (Zhou et al., 2024d)
W.M. KiGRAS (Zhao et al., 2024) / SMART (Wu et al., 2024b)
MUVO (Bogdoll et al., 2023)
Vista (Gao et al., 2024) / GAIA-1 (Hu et al., 2023a)
SIM. Waymax (Gulino et al., 2024)
SUMO (Krajzewicz et al., 2012) / LimSim (Wenl et al., 2023)
CARLA (Dosovitskiy et al., 2017)
STRIVE (Rempe et al., 2022)
MetaDrive (Li et al., 2022a)
Unisim (Yang et al., 2023b) / OAsim (Yan et al., 2024a)
NeuroNCAP (Ljungbergh et al., 2024)
DriveArena (Yang et al., 2024)
HUGSIM (Zhou et al., 2024a)
@@ 2 Related work
Autonomous driving. Recent research in autonomous driving have shifted from addressing individual tasks to exploring end-to-end planning, enabling the progressive execution from perception to ego-planning within a unified framework. Early works (Casas et al., 2021; Hu et al., 2022; Chitta et al., 2023) implement this by leveraging simplified auxiliary tasks, which limits the final performance. In contrast, UniAD (Hu et al., 2023b) and VAD (Jiang et al., 2023) have advanced this paradigm by integrating a broader spectrum of driving tasks, achieving notable progress in planning tasks by producing explicit intermediate results. Additionally, recent approaches present sparse representations (Zhang et al., 2024a; Sun et al., 2024), employ lightweight diffusion-based generation (Liao et al., 2025) to efficiently facilitate end-to-end planning. For enabling these driving methods to be evaluated extensively and authentically, we provide a reliable testing platform here.
@@
Sensor simulation. The geometric reconstruction of extensive urban spaces, such as streets and cities, has played a pivotal role in autonomous driving (Sun et al., 2020; Caesar et al., 2020). Recently, Gaussian splatting (Kerbl et al., 2023) based methods have been introduced to model dynamic urban scenes. PVG (Chen et al., 2023) incorporates periodic vibration at each Gaussian primitive to represent static and dynamic objects uniformly. At the same time, explicitly decomposing scenes into independent entities has become common practice, as seen in works such as StreetGaussians (Yan et al., 2024b), DrivingGaussians (Zhou et al., 2024c), HUGS (Zhou et al., 2024b) and OmniRe (Chen et al., 2024). Recently, LiDAR simulation studies have focused on using real data for improved realism. For instance, LiDARsim (Manivasagam et al., 2020) and PCGen (Li et al., 2023) use multi-step, data-driven methods to simulate point clouds from real data. Additionally, works such as (Tao et al., 2023; Zhang et al., 2024b; Zheng et al., 2024; Xue et al., 2024; Tao et al., 2024; Wu et al., 2024a) leverage NeRF (Mildenhall et al., 2020) for scene reconstruction and LiDAR simulation. Recent works (Jiang et al., 2025; Zhou et al., 2025) introduce Gaussian splatting (Kerbl et al., 2023) into the LiDAR reconstruction task, achieving improved reconstruction quality and rendering speed. However, sensor-acquired data is prone to occlusions, leading to information loss within scenes. Additionally, reconstructed dynamic objects can distort when viewed from alternative perspectives, limiting scene editability and novel view synthesis. To address these challenges, we propose modeling the foreground and background separately allowing them to be composed in a flexible manner, resulting in a comprehensive driving simulation platform.
@@
Closed-loop simulation in realistic settings. Closed-loop simulation (Gulino et al., 2024; H. Caesar, 2021; Dosovitskiy et al., 2017; Li et al., 2022a) is crucial for the evaluation and deployment of AD planning systems, collecting sufficient driving statistics and improving the emergency response capability of driving agents. To make simulators more realistic, recent research has explored using existing real-world driving datasets. Bench2Drive (Jia et al., 2024) improves the CARLA benchmark by reconstructing the data format to be more aligned with the nuScenes dataset, bridging the gap between reinforcement learning planners and end-to-end planners. Navsim (Dauner et al., 2024) imitates closed-loop evaluation to adjust the nuPlan benchmark, introducing more comprehensive and practical metrics. With the emergence of novel 3D rendering and generation research, some works integrate these techniques into existing datasets, better simulating and constructing richer and more diverse scenarios. NeuroNCAP (Ljungbergh et al., 2024) utilizes NeRF (Mildenhall et al., 2020) to render novel surrounding views and creates collision scenarios to enhance measurement. Relying on powerful generative models and extensive driving data, DriveArena (Yang et al., 2024) is capable to yield controllable and abundant real-world scenarios and realize closed-loop simulation. However, these simulators all fail to meet the fundamental criteria as mentioned earlier. This motivates the introduction of our RealEngine, a more reliable and powerful benchmark for assessing the real-world performance of driving agents.
@@ 3 Method
We present RealEngine, a reliable and comprehensive autonomous driving simulation platform capable of rendering realistic multimodal sensor data efficiently and supporting closed-loop simulation. It is composed of: (i) Simulator infrastructure including background scene and traffic participants as well as their compositions; (ii) Closed-loop driving simulation; (iii) Assessment of driving agents. The working mechanism of RealEngine is depicted in Figure˜1.
@@ 3.1 Simulator infrastructure
The advancement of recent generative models has been recently exploited for closed-loop driving simulation. For example, DriveArena (Yang et al., 2024) streamlines a couple of generative functions, such as interactive traffic flows and novel scene synthesis, allowing to simulate virtual driving scenes. However, this approach suffers from several key drawbacks: (1) Need for large manually labelled training data for video generative model optimization; (2) Substantial running overhead, even without support to multimodal sensor data such as LiDAR point clouds (3) Inability to support multi-agent co-operation and interaction. (4) Dependence on a pre-defined high-definite map; (5) Low controllability on the traffic scenario including background scene and traffic participants, consequently causing frequent situational inconsistency over space and time; (6) Low spatiotemporal realism. While some of the above issues, such as realistic scene rendering, more flexible traffic scenario editing, and no need for training data collection, can be addressed by NeuroNCAP (Ljungbergh et al., 2024), it narrowly focuses on safety test of driving agents. Further, it cannot well generalize to free form trajectories otherwise the novel view synthesis will degraded substantially.
@@ 3.1.1 Background scene models
To reconstruct realistic background environments, we adopt StreetGaussians (Yan et al., 2024b) for camera images and GS-LiDAR (Jiang et al., 2025) for LiDAR point clouds, chosen for their high rendering efficiency and cross-modal fidelity. Importantly, RealEngine maintains flexibility—alternative methods such as PVG (Chen et al., 2023), OmniRe (Chen et al., 2024), and LiDAR-RT (Zhou et al., 2025) can be seamlessly incorporated into our system as modular components.
Pose calibrations in nuPlan (H. Caesar, 2021) is relatively coarse, posing challenges for accurate scene reconstruction. A widely used conventional method for camera pose correction (Yan et al., 2024b; Chen et al., 2024; Hess et al., 2024) involves learning a trainable correction matrix that adjusts the camera pose automatically during training. However, this method is only effective for small pose deviations and fails to converge on nuPlan, leading to suboptimal results. To address this issue, we propose a pose correction method based on LiDAR point cloud registration.
As the geometric information of the background scene remains fixed in the world coordinate system, we can align the ego vehicle’s poses across different frames by registering LiDAR points transformed to the world. We remove dynamic vehicles from the LiDAR data based on annotation information and filter out ground points which vary significantly across frames. All remaining point clouds are transformed into the world coordinate system and truncated within a predefined region to ensure consistency across frames. We select the central LiDAR frame ${\bm{P}}_{\mathrm{refer}}$ as the reference frame and apply a learnable correction matrix $\bm{M}$} to transform the LiDAR frames $\bm{P}$}. The transformed LiDAR frames are then compared with the reference frame using the Chamfer Distance (C-D) (Fan et al., 2017) as the loss function:
@@
Additionally, we leverage learnable image exposure transformations to handle cross-camera appearance variations, and utilizes video generative prior (Yang et al., 2025; Yu et al., 2024a) to optimize the scene model across multiple trajectories in Section˜A.1.2. For further details on camera image and LiDAR point cloud reconstruction, please refer to Section˜A.1. By leveraging these advanced reconstruction techniques, we achieve precise multi-modal background reconstruction with high rendering efficiency, enabling flexible scene editing and realistic simulation of autonomous driving agents.
@@ 3.1.2 Traffic participants models
Traffic participants (e.g., vehicles, bicycles) are essential for realistic driving scenarios. To enable high-fidelity and diverse simulations, we curate a comprehensive set of 3D participant models. Conventional approaches (Yan et al., 2024b; Chen et al., 2024; Zhou et al., 2024a) insert reconstructed Gaussian splatting primitives into the background scene according to predefined trajectories. While providing high-quality rendering from training viewpoints, they degrade significantly when rendering resolution, viewpoint, or object distance changes (Yan et al., 2024c; Yu et al., 2024b; Song et al., 2024). This degradation causes visual artifacts, especially in close-range interactions (e.g., collisions), and also introduces inconsistencies in lighting and shadows between foreground objects and the background scene.
To address these issues, RealEngine represents traffic participants using 3D meshes, ensuring consistent geometry across all viewpoints and distances. To achieve seamless integration, we employ a diffusion model to guide learning of scene-consistent lighting and shading, as further detailed in Section˜3.1.3. Meshes are rendered using ray tracing to produce both RGB images and depth maps, enabling accurate occlusion reasoning between participants and scenes.
The 3D meshes are sourced from two complementary pipelines: (i) High-quality meshes manually created for key traffic objects. (ii) Meshes reconstructed from real-world datasets, including 360-degree images from CO3D (Reizenstein et al., 2021) and 3DRealCar (Du et al., 2024). These are processed via 3D Gaussian Splatting (3DGS) for reconstruction, followed by mesh extraction for flexible use in scene composition.